In a simple model of Earth's oceans, the upper layer of water that is well mixed by waves, wind, and heating by sunlight is about 500 m deep. This layer sits on a deeper layer of colder water beneath it. The area of the upper layer is about 3.60e+08 km2 and its average temperature is currently about 17 °C. Suppose that the temperature of this upper layer is increased to 19.2 °C. (a) By how much will the volume of this layer increase?
Answer #1

we know

dV = \alpha*V*dT

where dV is the change in volume; V is the initial volume = 500*3.6*108*106 m3 = 1.8*1017 m3; \alphais volume expansion coeffecient = 0.000214 per oC; dT is the change in temperature, dT = 19.2-17 = 2.2 oC

putting all the values

dV = 0.000214*1.8*1017*2.2

dV = 8.4744=1013 m3